a person standing on vertical cliff at height of 180 m above a lake,wantsto jump into the lake but notiecs a rock at the surface level with its furthest edge a distance s from the shore the person realises that with the running start it will be possile to just clear the rock.so the person steps back from the edge a distance d=125/3 and starting freom rest runs at an accelearation that varries in time according to a(x)=2t and then leaves the cliff horizontally.the person just clears the rock.find k if s=k50(take g=10cm/sec(^2) )
1
Expert's answer
2021-07-29T10:45:01-0400

"\text{Consider the movement of a person along a cliff:}"

"a(x)=2t"

"a(x) = v'(t)"

"v(t)= t^2+C_1"

"v(0) = 0\implies C_1=0"

"v(t)= t^2"

"s(t) = v'(t)"

"s(t) = \frac{t^3}{3}+C_2"

"s(0) = 0\implies C_2=0"

"s(t) = \frac{t^3}{3}"

"d = \frac{125}{3}"

"s(t_1) =\frac{125}{3}"

"\frac{t_1^3}{3}=\frac{125}{3}"

"t_1=5s;v_x(t_1)=t_1^2=25\frac{m}{s}"

"v_x - \text{speed of a person at the edge of a cliff}"

"\text{Consider a flight from the edge of the cliff}"

"\text{to the surface of the lake:}"

"\text{vertical movement}"

"h = 180m"

"h =\frac{gt_2^2}{2}"

"t_2=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2*180}{10}}=6s"

"\text{horizontal movement}"

"s = v_xt_2"

"s=50k"

"k= \frac{v_xt_2}{50}= \frac{25*6}{50}=3"

"\text{Answer: }k =3"